OpenStudy (anonymous):

whats the square root of 200 ?

OpenStudy (anonymous):

14.1421

OpenStudy (anonymous):

14.142

OpenStudy (michele_laino):

@shawntaria please remember that 200 = 2 * 10^2, then....

OpenStudy (anonymous):

\[10\sqrt{2}\]

OpenStudy (anonymous):

14.142

OpenStudy (anonymous):

\[10\sqrt{2}\]